DETECTION OF THE MAGNETIC DIPOLE-ALLOWED ORIGIN BAND OF THE Ã←X̃ TRANSITION OF THE NITRATE RADICAL NO_{3}

\underline{KANA TAKEMATSU}, NATHAN EDDINGSAAS, and MITCHIO OKUMURA, Arthur Amos Noyes Laboratory of Chemical Physics, California Institute of Technology, Pasadena, California 91125; JOHN STANTON, Department of Chemistry, University of Texas at Austin, Austin, TX 78712.

Despite its simplicity and atmospheric significance, the lowest electronic states of NO_{3} remain poorly understood. The Ã^{2}E′′ state is Jahn-Teller (JT) active and is pseudo-JT coupled to the ground X̃^{2}A_{2}′ and excited B̃^{2}E′′ states. While the band origin 0_{0}^{0} of the Ã←X̃ transition is purely electric forbidden, the cavity ringdown spectrum of the vibronicallyallowed Ã←X̃ bands has shown rotationally-resolvable structures. We have detected the magnetic dipole-allowed origin (∼7060 cm^{−1}) via CRD spectroscopy. Analysis of the band contour along with its implications of the JT/Pseudo-JT couplings in the dark Ã state of NO_{3} will be discussed.